OpenStudy (anonymous):

factor the expression: 2p^2-15p-8

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-16)(x+1) since you have a coefficient in front you need to divide so you get (2x- 16)(x+1/2)

OpenStudy (anonymous):

(2p+1)(p-8)

OpenStudy (anonymous):

@salma1996 how did you factor it?

OpenStudy (anonymous):

@kaisan I calculated the discriminant which is 15^2+4*8*2=17 and then I calculated the roots, the first one is (15+17)/2=8 and the second one is (15-17)/2=-1/2 and afterhat you can factor it like this 2(p+1/2)(p-8)=(2p+1)(p-8). I don't know if you've studied this but it's how you solve quadric equations.

OpenStudy (anonymous):

oh ok thanks ^-^